The siggraph Document Class Users’ Guide^{∗}

Daniel Kartch^{†}

1996/01/17

Abstract

This document class modifies the standard article class to conform to the specifications for papers published in the proceedings of the annual ACM Siggraph conference. It sets all the necessary layout parameters and handles the differences in format between a paper being submitted for blind review and a camera ready copy of an accepted paper. Several additional features are also provided.

DISCLAIMER: This class is distributed in the hope that it will be useful, and to promote compatibility of documents created at different sites. However it is provided WITHOUT ANY WARRANTY; without even the implied warranty of MER-CHANTABILITY or FITNESS FOR A PARTICULAR PUR-POSE. Please see the siggraph.dtx source file for copyright and distribution information.

1 Introduction

This document describes the siggraph document class, for use with L^{A}TEX 2ε. It is intended as a replacement for the various versions of the old siggraph.sty file that have been floating around for years. It loads in the article class and modifies several parameters to comply with the layout specifications for a paper submitted to the ACM Siggraph conference. Currently, these specifications are:

• \frac{3}{4} inch margins on top and sides, 1 inch margin on bottom

• two column mode with 2 pica column separation

• space for copyright is 1.5 inch

• 9 point type on 10 point baselines

• no page numbers

Also set by the class but not part of the siggraph specifications are:

• titles and section headings are bold sans serif

• 1 em paragraph indentation

• flushbottom mode

∗This file has version number v1.1.1 dated 1996/01/17. The most recent version can be found at http://www.graphics.cornell.edu/~dan/texstuff/siggraph/

†dan@graphics.cornell.edu

The class also allows the user to specify whether the paper is being submitted for blind review or as a final, camera-ready document. When review mode is specified, it will automatically generate a cover sheet for the paper, modify the title block on the first page to protect anonymity, and (optionally) switch to a more widely spaced layout. Mechanisms are also provided to allow the user to introduce text conditional on the chosen submission mode.

2 Usage

To use this class, specify

\documentclass{siggraph}

at the beginning of your document. The following options may be included in the

\documentclass command:
cameraready  • cameraready, preprint, review, or widereview
preprint
review
widereview

These specify the submission mode for the paper. cameraready, the default, sets the layout parameters as described in section 1. The preprint option adds page numbers and prints the message “To appear in the SIGGRAPH conference proceedings” in the top margin.

\preprinttext This message can be changed with the \preprinttext{⟨message⟩} command. The review option uses the same layout as cameraready but adds page numbers and modifies the title block as described in the next section. The widereview option is for people who prefer review papers that have a more widely spaced layout. It uses one-column mode with 12 point, double-spaced type. Other effects of these options will be described later.

onecolumn  • onecolumn or twocolumn
twocolumn
• singlespace or doublespace
singlespace
doublespace

• 9pt, 10pt, 11pt, or 12pt
9pt
10pt
11pt
12pt

These three sets of options will override the default settings for the number of columns, line spacing, and point size determined by the submission mode. Overriding the settings when in camera-ready mode, while allowed, is a violation of the submission specifications and will produce a warning message.

version1996  • version1996 or version1994

version1994 The submission specifications can change from year to year, and I will try to keep this class updated to match them. However, since many papers are painstakingly typeset to fit the allotted pages exactly, small changes in the format can radically change their appearance. This class therefore includes a specification date to allow old documents to maintain their appearance even if the specs change. Although the most recent version is always the default, I recommend including it in the options list anyways, so that you won’t have to think about it several years down the road. (The only difference between the 94 and 96 versions is that the space for the copyright is now a half inch larger.)

• Additionally, any other options supported by the article class may be used, except for landscape and titlepage, which are disabled.

3 Title Block, Abstract, and Cover Page

3.1 Data fields

In addition to the \title and \author commands provided by the article class, several other pieces of information can be set. Each of these commands takes a single argument, the value to which the corresponding field should be set.

\affiliation  • \affiliation

If all the authors share an affiliation, it can be set with this command rather than individually within the author command. The affiliation will appear centered below the author list.

\category  • \category and \format

\format The paper category (research, systems, ...) and format (print, video).

\contactname  • \contactname, \contactaddress, \contactphone, \contactfax, \contactemail
\contactaddress
\contactphone
\contactfax
\contactemail

These should be used to set the information about the person to be contacted about the paper.

     • \passport
\passport

The author who will receive compensation if the paper is accepted. As of 1996, this is no longer requested when submitting papers for review and the command has been disabled.

\keywords  • \keywords

A list of keywords to be printed on the cover page

\estpages  • \estpages

The estimated number of pages that the paper would require in the conference proceedings.

3.2 Title block

\maketitle The title block is generated with the normal L^{A}TEX \maketitle command. For a camera ready document it will contain:

• Title

• Authors

• Affiliation (if any)

For a review paper it will instead contain:

• Title

• Paper category

The review options will also suppress printing of any information specified with the \thanks command.

Some authors like to include a pretty picture of their results between the title block and the start of the text to act as a teaser. Because of the way L^{A}TEX’s float placement scheme works, this seemingly simple task is apparently impossible to do in a two column document without making the picture part of the \maketitle command. Therefore, that’s exactly what I’ve done. The class pro-

\teaser vides a \teaser{⟨figure⟩} command, whose contents will appear as a figure centered below the title block. Theoretically, anything that can appear in a figure environment can appear here. However, the current implementation is a bit of a kludge, so some things might not work. I’ve already been told that it isn’t compatible with the float package. I’ll see what I can do to improve it when I get the chance.

The space between the title block text and the rest of the document, and between the teaser (if any) and the start of the main text are controlled by the

\titlespace length parameters \titlespace and \teaserspace. They both default to \frac{1}{4} inch, \teaserspace but can be changed with the normal L^{A}TEX length setting commands.

3.3 Abstract

The abstract should be specified with the abstract environment. In addition to appearing in the paper itself, its contents will stored and will appear on the cover page (in review mode).

3.4 Cover page

When the review or widereview option is specified, a cover sheet will automat-

\suppresscover ically be generated at the end of the document. Placing the \suppresscover command in the preamble will prevent this. The cover sheet will contain:

• Title

• Authors

• Affiliation (if any)

• Paper category

• Contact information

• A copy of the abstract

• List of keywords

• Estimated number of pages

3.5 Obsolete command

In previous versions of this class, it was necessary to use a command called

\acmopening  \acmopening in place of the \maketitle command and abstract environment

in order to get the abstract to appear on the cover page as well as in the text. This is no longer necessary, and users should switch back to the normal commands.

4 Other Features

4.1 Copyright notice

\copyrightspace As with the old siggraph.sty, this class provides a \copyrightspace command

to leave space at the bottom of the first column for the copyright notice. This command should be placed after the last footnote that will appear in the first column. As of yet, there is no way to have this done automatically, but I am looking into it.

4.2 Title and section heading font

By default, this class uses bold sans-serif for the title and section heading font.

\sectionfont This behavior can be altered with the \sectionfont{⟨fontcommands⟩} command. ⟨fontcommands⟩ should be one or more font changing commands, for instance

“\rmfamily\scshape” or “\fontfamily{pcr}\selectfont”.

4.3 Conditional text

\ifcamera The class also provides the following four commands for writing text conditional

\ifreview
\ifcameraelse
\ifreviewelse

on whether camera-ready or review mode has been chosen:

\ifcamera{⟨thenclause⟩}

\ifreview{⟨thenclause⟩}

\ifcameraelse{⟨thenclause⟩}{⟨elseclause⟩}

\ifreviewelse{⟨thenclause⟩}{⟨elseclause⟩}

The first two will generate the ⟨thenclause⟩ if camera-ready (resp. review) mode is chosen, otherwise they resolve to a null string. The second two are the same as the first two but generate the ⟨elseclause⟩ instead of a null string if the condition is not met.

These can be used for writing the paper so as to protect anonymity for a submission for review, without having to worry about going back over it months later to put the identifying information back in if the paper is accepted. For example:

     In \cite{FooBar1994}, \ifcameraelse{we}{Foo and Bar} showed that ...

or

     as illustrated in the following \ifcameraelse{three}{two} images:

[include generic image]

\ifcamera{ [include our logo] }

[include another generic image]

4.4 Configuration file

Individual users or sites may wish to modify the behavior of this class (for instance, to use a particular set of fonts.) Rather than having people hack the .cls file, which can lead to various incompatible versions at different sites, the class provides for a configuration file. After the class has finished loading, it will check to see if a siggraph.cfg file exists in the TEX search path. If so, it will be input. Any user– or site–specific modifications should be placed in this file. (See the copyright notice in the siggraph.dtx source file for information on distribution of configuration files.)